A ball is thrown horizontally from a height of 20 meter and hits the ground with a speed that is three times its initial speed what is the initial speed?
1
Expert's answer
2019-12-09T09:03:25-0500

The law of conservation of energy gives


"mgh+\frac{mv_i^2}{2}=\frac{mv_f^2}{2}"

"mgh+\frac{mv_i^2}{2}=\frac{m(3v_i)^2}{2}"

So


"gh=4v_i^2"

"v_i=\frac{\sqrt{gh}}{2}=\frac{\sqrt{9.8\:\rm m/s^2\times 20\:\rm m}}{2}=7\:\rm m/s"